2
QuizMultiple Choice

The Multiplication Rule for Independent Events

Question 2 • AZ-Common Core Algebra II A

At the beginning of the semester, a professor tells students that if they study for the tests, then there is a 55% chance they will get a B or higher on the tests. If they do not study, there is a 20% chance that they will get a B or higher on the tests. The professor knows from prior surveys that 60% of students study for the tests. What is the probability that a randomly selected student studies for a test and gets a B or higher?

Answer
A
0.08
B
0.11
C
0.33
D
0.55

Explanation

The probability that a student both studies and earns a B or higher is the probability of studying multiplied by the probability of earning a B or higher given that the student studies. Thus, 0.60 × 0.55 = 0.33. The correct choice is 0.33.

AI-written and checked against the verified answer.

Previous
Next